AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 183     Entry time: Tue Feb 1 14:43:20 2022
Author: JM, TD 
Subject: Detector Bias, FEEs 

Detector is not biasing as it has in previous tests. Last bias 15/1/22 see previous elogs, max leakage current ~10 uA at 150V

Detector bias increases well above 10uA for voltages as low as 30-40V. Optical windows covered, Y09 valves closed, large ion pump OFF and both ion gauges OFF produces similar results but with a slightly lower leakage current.

Grounding and pulser cables removed from FEEs and jumper attached to nn bias adaptor card - no change. Bias module swapped out - no effect. NIM rack shows expected voltages (6,12,24) on multimeter.

 

AIDA03 - disconnected. Behaves as expected. Get waveforms and pulser peak. Pulser peak FWHM ~ 70 channels, noise transients observed.

AIDA02 - Not producing ADC data

AIDA01,04,05,06,07,08 - Very high rates - noisy

 

Attachment 1: 22-02-01_15-03-56_aida03pulserADCspectra.png  206 kB  | Hide | Hide all
22-02-01_15-03-56_aida03pulserADCspectra.png
Attachment 2: 2022-02-01_15-14-07_aida03_pulser_DAQserver.png  202 kB  | Hide | Hide all
2022-02-01_15-14-07_aida03_pulser_DAQserver.png
Attachment 3: 2022-02-01_15-22-23aida03_waveforms_pulser.png  334 kB  | Hide | Hide all
2022-02-01_15-22-23aida03_waveforms_pulser.png
Attachment 4: 2022-02-01_15-28-45aida03_waveforms_pulser2.png  241 kB  | Hide | Hide all
2022-02-01_15-28-45aida03_waveforms_pulser2.png
Attachment 5: 2022-02-01_15-54-57rates_aidaALL.png  108 kB  | Hide | Hide all
2022-02-01_15-54-57rates_aidaALL.png
ELOG V3.1.4-unknown